# Look-Ahead Reasoning on Learning Platforms

### v1

@@ A.6 Proof of Theorem˜6
A.7 Proof of ˜7
@@
A.8 Proof of ˜8
The proof is the same as the proof of ˜7 up to a use of the envelope theorem. For completeness, we restate the proof here and highlight the use of the envelope theorem. For notational simplicity, we abbreviate $h^{\sharp}_{\alpha}$ as $h$. Consider the derivative of $U_{\alpha}^{*}$ with respect to $\alpha$,
@@ A.8 Proof of ˜8
A.9 Proof of ˜9
By Lemma˜11, the sensitivity of the mixture distribution can be computed as
@@ Appendix B Example to illustrate Theorem 6
In this example, ˜1 is satisfied. Moreover, $\Phi^{2}>0$ since $\Phi=-2r^{2}$ where $r:=\|\mathbf{p}_{3}\|_{2}$. The benefit of coordination $\mathrm{B}=U(h^{\sharp})-U(h^{*})=\frac{3}{4}r^{2}$ is strictly positive as long as the anchor points are appropriately spaced. One can check that $U$ is $2r^{2}$-strongly concave, and Theorem˜6 can be verified since $\mathrm{B}\leq\frac{\Phi^{2}}{2\gamma}=r^{2}$ which is tight up to a factor 1/4 coming from the slack in the strong concavity assumption.

### v2

@@ A.5 Proof of Theorem˜5
A.6 Proof of ˜5
For notational simplicity, we set $f(\eta,\theta):=\mathbb{E}_{z\sim\mathcal{D}_{h(\eta)}}\left[u(z;\theta)\right]$ and $g(\eta,\theta):=\mathbb{E}_{z\sim\mathcal{D}_{h(\eta)}}\left[\ell(z;\theta)\right]$. Recall that
$U(h(\eta))=f(\eta,\mathcal{A}(\mathcal{D}_{h(\eta)}))=\mathbb{E}_{z\sim\mathcal{D}_{h(\eta)}}\left[u(z,\mathcal{A}(\mathcal{D}_{h(\eta)}))\right],$
where the first argument in $f$ only applies in distribution that is taken against and the second argument is corresponding to the second argument of $u$. Then, by the implicit function theorem we have
$\nabla_{\eta}U(h(\eta))=\nabla_{\eta}f(\eta,\mathcal{A}(\mathcal{D}_{h(\eta)}))=\nabla_{1}f(\eta,\mathcal{A}(\mathcal{D}_{h(\eta)}))-\nabla^{2}_{1,2}g(\eta,\mathcal{A}(\mathcal{D}_{h(\eta)}))\left[\nabla^{2}_{2,2}g(\eta,\mathcal{A}(\mathcal{D}_{h(\eta)}))\right]^{-1}\nabla_{2}f(\eta,\mathcal{A}(\mathcal{D}_{h(\eta)})),$
where $\nabla_{1}$ and $\nabla_{2}$ denote the gradient operator on the first/second argument of the function. For the equilibrium of the selfish strategy $(h(\eta^{*}),\theta^{*})$, we must have $\nabla_{1}f(\eta,\mathcal{A}(\mathcal{D}_{h}))=0$. Let $\eta^{\sharp}:=\operatorname*{arg\,max}_{\eta}U(h(\eta))$, by the PL-inequality, we obtain that
$U\left(h(\eta^{\sharp})\right)-U\left(h(\eta^{*})\right)\leq\frac{1}{2\gamma}\cdot\left\|\nabla^{2}_{1,2}g(\eta^{*},\mathcal{A}(\mathcal{D}_{h(\eta^{*})}))\left[\nabla^{2}_{2,2}g(\eta^{*},\mathcal{A}(\mathcal{D}_{h(\eta^{*})}))\right]^{-1}\nabla_{2}f(\eta^{*},\mathcal{A}(\mathcal{D}_{h(\eta^{*})}))\right\|_{2}^{2}.$
A.7 Proof of ˜6
@@
A.8 Proof of ˜7
The proof is the same as the proof of ˜6 up to a use of the envelope theorem. For completeness, we restate the proof here and highlight the use of the envelope theorem. For notational simplicity, we abbreviate $h^{\sharp}_{\alpha}$ as $h$. Consider the derivative of $U_{\alpha}^{*}$ with respect to $\alpha$,
@@ A.8 Proof of ˜7
A.9 Proof of ˜8
By Lemma˜10, the sensitivity of the mixture distribution can be computed as
@@ Appendix B Example to illustrate Theorem 5
In this example, ˜1 is satisfied. Moreover, $\Phi^{2}>0$ since $\Phi=-2r^{2}$ where $r:=\|\mathbf{p}_{3}\|_{2}$. The benefit of coordination $\mathrm{B}=U(h^{\sharp})-U(h^{*})=\frac{3}{4}r^{2}$ is strictly positive as long as the anchor points are appropriately spaced. One can check that $U$ is $2r^{2}$-strongly concave, and Theorem˜5 can be verified since $\mathrm{B}\leq\frac{\Phi^{2}}{2\gamma}=r^{2}$ which is tight up to a factor 1/4 coming from the slack in the strong concavity assumption.